-10x+3=-3x+12-4x
I'm not sure where to start on this. I can usually (but definitely not always) get it if I can get started. Please help!!
The more help the better!!!

Answer by wuwei96815(245)

The easiest way to solve an equation like this is to get all like terms on one side of the equations.

-10x + 3 = -3x + 12 -4x
-10x +3x +4x = 12 - 3
-3x = 9
x = 9/-3 = -3

Substitute -3 into the original equation to check the answer.